OpenStudy (anonymous):

y varies directly as the square of x. when x=3, y=8. Find y when x=2

OpenStudy (anonymous):

sorry what;s the answer ?

OpenStudy (angela210793):

f(x) is \[y=\sqrt{x}\]

OpenStudy (angela210793):

Sorry I lost it :(

OpenStudy (anonymous):

no...it should be done in this way....

OpenStudy (anonymous):

\[y=k{x^2}\] when x=3, y=8 8=k*3^2 k=8/9 y=8x^2/9 when x=2 y=8*2^2/9=32/9